# Nested Sampling for Exploring Lennard - Jones Clusters

### v1

@@ Abstract
Lennard - Jones clusters, while an easy system, have a significant number of non equivalent configurations that increases rapidly with the number of atoms in the cluster. Here, we aim at determining the cluster partition function; we use the nested sampling algorithm, which transforms the multidimensional integral into a one-dimensional one, to perform this task. In particular, we use the nested_fit program, which implements slice sampling as search algorithm. We study here the 7-atoms and 36-atoms clusters and then look at the computational cost of nested_fit with and without parallelisation. We find that nested_fit is able to recover phase transitions and find different stable configurations of the cluster. Furthermore, the implementation of the slice sampling algorithm has a clear impact on the computational cost.
Keywords
@@ 1 Introduction
Lennard - Jones clusters are relevant models for rare gas atoms clusters [1] but the number of configurations increases exponentially with the number of degrees of freedom [2]. They have been widely studied using a variety of methods such as basin-hopping [3], Metropolis Monte-Carlo [4], Molecular Dynamics [5] or parallel tempering [6, 7].
In this work, we use the nested sampling algorithm [8] to study those clusters. This algorithm is mainly employed to analyse data in fields such as cosmology and astrophysics [9] but is applied more and more in materials science for the exploration of potential energy surfaces, including Lennard - Jones clusters [10]. Here, we utilise a different implementation of nested sampling than the one used in Ref. [10].
@@ 2 Nested sampling for Lennard - Jones clusters
For a cluster of $N$ atoms, the truncated and shifted Lennard - Jones potential takes the following form [11, 12]
$V=\sum_{1\leq i<j\leq N}\tilde{V}(r_{ij})$ (1)
with
$\tilde{V}(r)=\begin{cases}V_{LJ}(r)-V_{LJ}(r_{c})&r<r_{c},\\ 0&r\geq r_{c},\end{cases}$ (2)
and
$V_{LJ}(r)=4\epsilon\left(\left(\frac{r_{0}}{r}\right)^{12}-\left(\frac{r_{0}}{r}\right)^{6}\right).$ (3)
@@
[…]
@@
    This procedure is repeated until the current contribution is small compared to previous contributions: $\log(c_{i})-\log(c_{max})<\delta$ with $c_{max}=\max_{m\leq i}(c_{i})$ at an inverse temperature $\beta$ chosen by the user. Here, we use $\delta=-10$, a value that satisfies our requirements.
As the interaction potential does not depend on the temperature, we can compute the partition function at all temperatures by estimating pairs of ($E_{i}$,$w_{i}$) values during one single exploration of the potential [10] as
$Z_{c}(\beta)\approx\sum_{i}c_{i}.$ (6)
@@
$U=-\frac{\partial\log(Z)}{\partial\beta}=-\frac{\partial\log(Z_{k})}{\partial\beta}-\frac{\partial\log(Z_{c})}{\partial\beta}=\frac{3N}{2}k_{B}T-\frac{\partial\log(Z_{c})}{\partial\beta}$ (7)
@@
$C_{v}=\frac{\partial U}{\partial T}=\frac{3N}{2}k_{B}-\frac{\partial}{\partial T}\frac{\partial\log(Z_{c})}{\partial\beta}.$ (8)
In this paper, we use, in particular, the nested_fit program [14, 15, 16, 17]. In the next sections, we present some of the specific implementation of the nested sampling algorithm in nested_fit.
2.1 Slice sampling and slice sampling transformed
The method used to find the new point $\boldsymbol{x}_{new}$ in nested_fit is slice sampling [18, 19]. This method consists in uniformly choosing new exploration points on a slice of the volume defined by the constraint $V(\boldsymbol{x})<V(\boldsymbol{x}_{old})$. In one dimension, one of the live points is randomly chosen and a slice is built around it until the end points of the slice have an energy higher than $\boldsymbol{x}_{old}$ or are out of the sampling space. A point is then sampled from within the slice and accepted if it verifies the constraint and rejected otherwise. In that case, another point is sampled until an acceptable point is found. In a multidimensional setting, a change of coordinates is first performed to efficiently explore all the parameter space, even in presence of strong correlation. This transformation is done via the Cholesky transformation of the covariance matrix of the live points of the considered step to transform the points coordinates into new coordinates with dimensions $\sim\mathcal{O}(1)$ in all directions [19]. The one-dimensional algorithm is then applied recursively to the vectors of $n_{bases}$ randomly generated orthonormal bases (see Ref. [17]). Hence, $n_{bases}\times 3N$ steps are performed.
@@ 2.1 Slice sampling and slice sampling transformed
    First, the steps can be performed in the transformed space. This will be referred to as slice sampling transformed and was used in Ref. [17]. However, the sampled points need to be transformed back to the real space to compute the energy and check that the points are within the bounds. This brings a significant computational overhead to the computation.
@@
    Second, the steps can be performed in the real space and only the slices are chosen in the transformed space. This will be referred to as slice sampling and was used in Refs. [20, 19]. In that case, only the orthonormal bases need to be transformed to the real space.
@@ 2.2 Parallelisation
The nested sampling algorithm can be easily parallelised. One method that is used to make nested sampling parallel is, instead of searching and substituting one point per iteration, to search $r$ new points in parallel and then to replace $r$ points at once [21]: the $r$ points with the highest energy are removed and replaced by $r$ new points with an energy lower than the energies of all removed points. In that case, the compression factor is $\frac{K-r+1}{K}$ instead of $\frac{K}{K+1}$ (see Eq. (5)) [13]. However, this parallelisation increases the variance of the results [13]. A different approach is used in Polychord using MPI protocols [20, 19]: one primary process is in charge of the sequential replacing of the live points while all other secondary processes are in charge of finding new live points continuously, with the threshold criterion changed by the primary process. The sequential replacing is thus done by the primary process with the points that are found by the secondary processes. A point found for an iteration with a specific energy constraint is valid for a further iteration if it verifies the new constraint [19]. In this case, the compression factor is still $\frac{K}{K+1}$.
@@
  • If the energy is higher than $V_{max}$, the new point is rejected. In this case, $V_{max}$ is not updated.
@@ 3 Lennard - Jones clusters
We now study Lennard - Jones clusters of two sizes : 7 atoms and 36 atoms. In the case of the 7-atoms cluster, we compare our results with those found in Ref. [10]. In this work, we use reduced units i.e. the temperature is in units of $\frac{k_{B}T}{\epsilon}$ with $\epsilon=1$. In reduced units, the stopping temperature is going to be 0.010010.010.01 for the 7-atoms cluster and 0.00500050.0050.005 for the 36-atoms cluster. For the simulation, the atoms are confined in a cubic box of side $L$. To fix the size of the box, we will look at the value of the density $\rho$ in unit of $r_{0}^{-3}$: $\rho=\frac{N}{(L/r_{0})^{3}}r_{0}^{-3}$. In all our cases, we take $L=6$ and use $r_{0}$ to tune the density to a given desired value. Within reduced units, the results are valid for all atomic species interacting via the Lennard - Jones potential: the specific ($r_{0}$,$\epsilon$) values, corresponding to the atomic species considered, need to be reintroduced instead of the reduced ones. Hence, the specific value of $r_{0}$ is not important. The density characterises the size of the space the particles evolve in, relative to the number of atoms $N$.
Here, we aim to compute the partition function and derivatives of the clusters, aiming at detecting the possible phase transitions. In a infinite system, a phase transition is indicated by a discontinuity in the heat capacity curve, either in the form of a singularity (delta function) or a critical exponent [22]. However, here we work in a finite system composed of $N$ particles in which case a phase transition is not shown by a discontinuity but rather by a peak in the heat capacity curve [23]. This peak can sometimes take the form of a shoulder.
@@
Figure 1: Heat capacity for the Lennard - Jones cluster with $7$ atoms. Comparison with the results in [10]. The configurations shown were obtained with XCrySDen [24]. We used $K=1000$ live points and $n_{bases}=5$. Eight runs were performed.
First, we consider a cluster made of $N=7$ atoms with a density of $\rho=9.27\times 10^{-3}r_{0}^{-3}$ (corresponding to $r_{0}=0.659$). The size of the cluster and the density are chosen in order to compare our results with those obtained in Ref. [10]. We use $K=1000$ live points repeating 8 times the procedure to estimate the uncertainties. Ref. [10] uses the pymatnest [10, 25, 26, 13] program with rejection Gibbs sampling to update iteratively the points — which moves one atom at a time — with $K=500$ live points. We both use a cutoff radius $r_{c}$ of $3r_{0}$. Both heat capacity curves obtained are represented in Figure 1 with examples of solid, liquid and gas configurations. We see that a good agreement with Ref. [10] is obtained. More importantly, nested_fit is able to recover both the evaporation and melting phase transitions. We also ran nested_fit with $K=500$ live points but obtained an evaporation peak that was slightly shifted towards higher temperatures compared to the curve from Ref. [10], indicating that there are too few points to obtain convergence. The difference in search method may explain why we need to double the points compared to Ref. [10]. Indeed, pymatnest was developed specifically for the exploration of energy surfaces while nested_fit has applications in both data analysis and materials science. The search method used by the latter thus have to work in both cases. Furthermore, in Ref. [10], the Lennard - Jones function was called $2\times 10^{7}$ times while here it was called $5.4\times 10^{7}$ times. There are two main reasons for this increase: first, we use twice as many live points, which should more or less double the number of calls, and second, in slice sampling, the function is called not only to move the points but also to extend the slice.
@@
Figure 2: Heat capacity for the Lennard - Jones cluster with $36$ atoms. The configurations shown were obtained with XCrySDen [24]. We used $K=70000$ live points and $n_{bases}=3$. Eight runs were performed. The full line is the mean obtained from the eight runs and the dashed curves correspond to the first standard deviation.
Second, we consider a cluster composed of $N=36$ atoms with a density of $\rho=1.79\times 10^{-2}r_{0}^{-3}$ (corresponding to $r_{0}=0.475$). This example is quite challenging as a low temperature peak is observed, associated to a Mackay - anti-Mackay phase transition [27, 28]. The heat capacity obtained with $K=70000$ is shown in Figure 2. As for the previous case, the calculation is repeated 8 times for the estimation of the uncertainties. Again, we can see the presence of a peak corresponding to evaporation and a shoulder corresponding to melting. However, in this case, there is also a peak at lower temperature (around $T=0.135$) that corresponds to a solid-solid phase transition. From the standard deviation of the heat capacity curve, we can see that the position of this peak is more unstable from one run to the other than the position of the evaporation peak. Indeed, this low temperature peak is harder to obtain — and was not seen in runs with fewer live points — as we need enough points to fall in each basin so that it is correctly explored. A variation in the proportion of samples in each basin could explain this difference between the runs. We found the positions of the peaks to be similar when more live points - up to $K=110000$ were used, indicating that $K=70000$ live points is enough to attain convergence. The two solid configurations are shown in Figure 2. The low temperature peak was also found in [10] at $T=0.145$ which is in the interval [0.1,0.15]delimited-[]01.015[0.1,0.15][ 0.1,0.15 ] given by our eight runs.

### v2

# Lennard-Jones; Clusters; Nested sampling

@@ Abstract
Lennard-Jones clusters, while an easy system, have a significant number of non equivalent configurations that increases rapidly with the number of atoms in the cluster. Here, we aim at determining the cluster partition function; we use the nested sampling algorithm, which transforms the multidimensional integral into a one-dimensional one, to perform this task. In particular, we use the nested_fit program, which implements slice sampling as search algorithm. We study here the 7-atom and 36-atom clusters to benchmark nested_fit for the exploration of potential energy surfaces. We find that nested_fit is able to recover phase transitions and find different stable configurations of the cluster. Furthermore, the implementation of the slice sampling algorithm has a clear impact on the computational cost.
keywords:
@@ 1 Introduction
Lennard-Jones clusters are relevant models for rare gas atoms clusters ashcroft_solid_1976 but the number of configurations increases exponentially with the number of degrees of freedom doye_evolution_1999. They have been widely studied using a variety of methods such as basin-hopping wales_global_1997, Metropolis Monte-Carlo quirke_melting_1984, Molecular Dynamics beck_interplay_1988 or parallel tempering neirotti_phase_2000; calvo_phase_2000.
In this work, we use the nested sampling algorithm skilling_nested_2004 to study those clusters. This algorithm is mainly employed to analyse data in fields such as cosmology and astrophysics ashton_nested_2022 but is applied more and more in materials science for the exploration of potential energy surfaces, including Lennard-Jones clusters partay_efficient_2010. Here, we utilise nested_fit trassinelli_nested_fit_2019; trassinelli_bayesian_2017; trassinelli_mean_2020; maillard_assessing_2023, which has a different implementation of nested sampling than the one used in Ref. partay_efficient_2010. Indeed, we want to benchmark nested_fit for the exploration of potential energy surfaces for classical and, in the future, quantum systems.
@@ 2 Nested sampling for Lennard-Jones clusters
For a cluster of $N$ atoms, the truncated and shifted Lennard-Jones potential takes the following form shi_histogram_2001; partay_polytypism_2017
$V=\sum_{\begin{subarray}{c}1\leq k<j\leq N\\ r_{i,kj}<r_{c}\end{subarray}}4\epsilon\left(\left(\frac{r_{0}}{r_{ij}}\right)^{12}-\left(\frac{r_{0}}{r_{ij}}\right)^{6}-\left(\frac{r_{0}}{r_{c}}\right)^{12}+\left(\frac{r_{0}}{r_{c}}\right)^{6}\right).$
@@
$\begin{split}Z(\beta)&=\frac{1}{N!~h^{3N}}\int\exp\left(-\beta\left(\frac{1}{2}\sum_{i=1}^{3N}\frac{p_{i}^{2}}{m}+V(\boldsymbol{x})\right)\right)d\boldsymbol{x}d\boldsymbol{p}\\ &=\frac{1}{N!~h^{3N}}\int\exp\left(-\beta\left(\frac{1}{2}\sum_{i=1}^{3N}\frac{p_{i}^{2}}{m}\right)\right)d\boldsymbol{p}\int\exp(-\beta(V(\boldsymbol{x}))d\boldsymbol{x}=Z_{k}*Z_{c},\end{split}$
where $\boldsymbol{p}=(p_{1},...,p_{3N})$ is the vector of momenta, $\boldsymbol{x}$ the vector of all atomic positions, $\beta=\frac{1}{k_{B}T}$ the inverse temperature and $h$ the Planck constant. The separation in the equation above is possible as the potential $V$ only depends on the atom positions and not on the momenta. The term $Z_{k}=\frac{1}{N!~h^{3N}}\int\exp\left(-\beta\left(\frac{1}{2}\sum_{i=1}^{3N}\frac{p_{i}^{2}}{m}\right)\right)d\boldsymbol{p}$ is easy to compute and gives $(\sqrt{2\pi m/(\beta h^{2})})^{3N}/N!$. The term $Z_{c}=\int\exp(-\beta(V(\boldsymbol{x}))d\boldsymbol{x}$ can be rewritten as $Z_{c}=\int\rho(E)\exp(-\beta E)dE$ where $\rho(E)$ is the density of states partay_nested_2021 and does not in general have an analytical form. We compute $Z_{c}$ by using a Monte Carlo sampling and, more precisely, the nested sampling algorithm which works in the following way skilling_nested_2004; partay_efficient_2010:
@@
    This procedure is repeated until the current contribution is small compared to previous contributions: $\log(c_{i})-\log(c_{max})<\delta$ with $c_{max}=\max_{m\leq i}(c_{m})$ at an inverse temperature $\beta$ chosen by the user. Here, we use $\delta=-10$, the value used in Ref. maillard_assessing_2023 that gave good results for the harmonic potential.
As the interaction potential does not depend on the temperature, we can compute the partition function at all temperatures by estimating pairs of ($E_{i}$$w_{i}$) values during one single exploration of the potential partay_efficient_2010 as
$Z_{c}(\beta)\approx\sum_{i}c_{i}.$
@@
$U=-\frac{\partial\log(Z)}{\partial\beta}=-\frac{\partial\log(Z_{k})}{\partial\beta}-\frac{\partial\log(Z_{c})}{\partial\beta}=\frac{3N}{2}k_{B}T-\frac{\partial\log(Z_{c})}{\partial\beta}$
@@
$C_{v}=\frac{\partial U}{\partial T}=\frac{3N}{2}k_{B}-\frac{\partial}{\partial T}\frac{\partial\log(Z_{c})}{\partial\beta}.$
In this paper, we use, in particular, the nested_fit program. In the next sections, we present some of the specific implementation of nested sampling in nested_fit.
2.1 Slice sampling real and slice sampling transformed
The method used to find the new point $\boldsymbol{x}_{new}$ in nested_fit is slice sampling neal_slice_2003; handley_polychord_2015-1. This method consists in uniformly choosing new exploration points on a slice of the volume defined by the constraint $V(\boldsymbol{x})<V(\boldsymbol{x}_{old})$. In one dimension, one of the live points is randomly chosen and a slice is built around it until the end points of the slice have an energy higher than $\boldsymbol{x}_{old}$ or are out of the sampling space. A point is then sampled from within the slice and accepted if it verifies the constraint and rejected otherwise. In that case, another point is sampled until an acceptable point is found. In a multidimensional setting, a change of coordinates is first performed to efficiently explore all the parameter space, even in presence of strong correlation. This transformation is done via the Cholesky transformation of the covariance matrix of the live points of the considered step to transform the points coordinates into new coordinates with dimensions $\sim\mathcal{O}(1)$ in all directions handley_polychord_2015-1. The one-dimensional algorithm is then applied recursively to the vectors of $n_{\textrm{bases}}$ randomly generated orthonormal bases (see Ref. maillard_assessing_2023). Hence, to compute an update candidate, one live point, corresponding to one $N$-atom cluster, is moved for $n_{\textrm{bases}}\times 3N$ steps, each one moving all atoms in the cluster. Increasing the value of $n_{\textrm{bases}}$ implies that there is a better decorrelation of samples but it also results in a higher computational cost.
@@ 2.1 Slice sampling real and slice sampling transformed
    First, the steps can be performed in the transformed space. This will be referred to as slice sampling transformed and was used in Ref. maillard_assessing_2023. However, the sampled points need to be transformed back to the real space to compute the energy and check that the points are within the bounds. This brings a significant computational overhead to the computation.
@@
    Second, the steps can be performed in the real space and only the slices are chosen in the transformed space. This will be referred to as slice sampling real and was used in Refs. handley_polychord_2015; handley_polychord_2015-1. In this case, only the orthonormal bases need to be transformed to the real space.
@@ 2.2 Parallelisation
The nested sampling algorithm can be easily parallelised. One method that is used to make nested sampling parallel is, instead of searching and substituting one point per iteration, to search $r$ new points in parallel and then to replace $r$ points at once henderson_parallelized_2014: the $r$ points with the highest energy are removed and replaced by $r$ new points with an energy lower than the energies of all removed points. In that case, the compression factor is $\frac{K-r+1}{K}$ instead of $\frac{K}{K+1}$ (see Eq. (1)) partay_nested_2021. However, this parallelisation increases the variance of the results partay_nested_2021. A different approach is used in Polychord using MPI protocols handley_polychord_2015; handley_polychord_2015-1: one primary process is in charge of the sequential replacing of the live points while all other secondary processes are in charge of finding new live points continuously, with the threshold criterion changed by the primary process. The sequential replacing is thus done by the primary process with the points that are found by the secondary processes. A point found for an iteration with a specific energy constraint is valid for a further iteration if it verifies the new constraint handley_polychord_2015-1. In this case, the compression factor is still $\frac{K}{K+1}$.
@@
  • If the energy is higher than $V_{max}$, the new point is rejected and $V_{max}$ is not updated.
@@ 3 Lennard-Jones clusters
We now study Lennard-Jones clusters of two sizes : 7 atoms and 36 atoms. In the case of the 7-atom cluster, we compare our results with those found in Ref. partay_efficient_2010. In this work, we use reduced units i.e. the temperature is in units of $\frac{k_{B}T}{\epsilon}$ with $\epsilon=1$. In reduced units, the stopping temperature is going to be $0.01$ for the 7-atom cluster and $0.005$ for the 36-atom cluster. For the simulation, the atoms are confined in a cubic box of side $L$. To fix the size of the box, we will look at the value of the density $\rho$ in unit of $r_{0}^{-3}$: $\rho=\frac{N}{(L/r_{0})^{3}}r_{0}^{-3}$. In all our cases, we take $L=6$ and use $r_{0}$ to tune the density to a given desired value. Within reduced units, the results are valid for all atomic species interacting via the Lennard-Jones potential: the specific ($r_{0}$$\epsilon$) values, corresponding to the atomic species considered, need to be reintroduced instead of the reduced ones. Hence, the specific value of $r_{0}$ is not important. The density characterises the size of the space the particles evolve in, relative to the number of atoms $N$. We use slice sampling real with the covariance matrix computed every $0.05K$ iterations and the procedure is repeated eight times to estimate the uncertainties.
Here, we aim to compute the partition function and derivatives of the clusters, aiming at detecting the possible phase transitions. In an infinite system, a phase transition is indicated by a discontinuity in the heat capacity curve, either in the form of a singularity (delta function) or a critical exponent huller_first_1994. However, here we work in a finite system composed of $N$ particles in which case a phase transition is not shown by a discontinuity but rather by a peak (or a shoulder) in the heat capacity curve medved_modeling_2017.
@@
Figure 1: (a) Heat capacity for the Lennard-Jones cluster with 77 atoms. Comparison with the results in Ref. partay_efficient_2010. The configurations shown were obtained with XCrySDen kokalj_xcrysdennew_1999. (b) Evolution of slice sampling’s acceptance ratio during one of the eight runs.
First, we consider a cluster made of $N=7$ atoms with a density of $\rho=9.27\times 10^{-3}r_{0}^{-3}$ (corresponding to $r_{0}=0.659$). The size of the cluster and the density are chosen in order to compare our results with those obtained in Ref. partay_efficient_2010. We use $K=1000$ live points and $n_{\textrm{bases}}=5$ so that the samples are decorrelated enough for the comparison. Ref. partay_efficient_2010, through the pymatnest partay_efficient_2010; baldock_determining_2016; baldock_constant-pressure_2017; partay_nested_2021 program, moves the atoms one at a time using rejection Gibbs sampling with $K=500$ live points. We both use a cutoff radius $r_{c}$ of $3r_{0}$. Both heat capacity curves obtained are represented in Figure 1 (a) with examples of solid, liquid and gas configurations. We see that a good agreement with Ref. partay_efficient_2010 is obtained. More importantly, nested_fit is able to recover both the evaporation and melting phase transitions. We also ran nested_fit with $K=500$ live points but obtained an evaporation peak that was slightly shifted towards higher temperatures compared to the curve from Ref. partay_efficient_2010, indicating that there are too few points to obtain convergence. The difference in search method may explain why we need to double the points compared to Ref. partay_efficient_2010. Indeed, pymatnest was developed specifically for the exploration of energy surfaces while nested_fit has applications in both data analysis and materials science. The search method used by the latter thus have to work in both cases. Furthermore, in Ref. partay_efficient_2010, the Lennard-Jones function was called $2\times 10^{7}$ times while here it was called $5.4\times 10^{7}$ times. There are two main reasons for this increase: first, we use twice as many live points, which should more or less double the number of calls, and second, in slice sampling, the function is called not only to move the points but also to extend the slice. In Figure 1 (b), we see that the acceptance ratio of a slice sampling search decreases when the volume decreases, showing that it is then more difficult to find new sampling points. The explorations were done on a 64 bits computer with 4 CPUs of frequency 3.3 GHz and took around one minute per exploration.
@@
Figure 2: Heat capacity for the Lennard-Jones cluster with 3636 atoms. The configurations shown were obtained with XCrySDen kokalj_xcrysdennew_1999. The full line is the mean obtained from the eight runs and the dashed curves correspond to the first standard deviation.
Second, we consider a cluster composed of $N=36$ atoms with a density of $\rho=1.79\times 10^{-2}r_{0}^{-3}$ (corresponding to $r_{0}=0.475$). This example is quite challenging as a low temperature peak is observed, associated to a Mackay - anti-Mackay phase transition mandelshtam_multiple_2006; frantsuzov_size-temperature_2005. Here, we use $n_{\textrm{bases}}=3$ as the energy function is more costly than in the previous case. The heat capacity obtained with $K=70000$ is shown in Figure 2. Again, we can see the presence of a peak corresponding to evaporation and a shoulder corresponding to melting. However, in this case, there is also a peak at lower temperature (around $T=0.135$) that corresponds to a solid-solid phase transition. From the standard deviation of the heat capacity curve, we can see that the position of this peak is more unstable from one run to the other than the position of the evaporation peak. Indeed, this low temperature peak is harder to obtain — and was not seen in runs with fewer live points — as we need enough points to fall in each basin so that it is correctly explored. A variation in the proportion of samples in each basin could explain this difference between the runs. We found the positions of the peaks to be similar when more live points — up to $K=110000$ — were used, indicating that $K=70000$ live points is enough to attain convergence. The two solid configurations are shown in Figure 2. The low temperature peak was also found in Refs. partay_efficient_2010; mandelshtam_multiple_2006; frantsuzov_size-temperature_2005 at around $T=0.145$ which is in the interval $[0.1,0.15]$ given by our eight runs. The study in Refs. partay_efficient_2010; mandelshtam_multiple_2006; frantsuzov_size-temperature_2005 were done at different densities, but this does not affect the position of the solid-solid transition peak. We also tried using $n_{\textrm{bases}}=1$ but we observed a shift of the evaporation peak: the number of bases used is probably too small to decorrelate the samples. The explorations were done on a 64 bits computer with 64 CPUs of frequency 2.0 GHz and took around 43 hours per exploration.